# Discrete time crystal for periodic-field sensing with quantum-enhanced precision

### v1

@@ IV DTC sensor
Fig. 2 (a) presents the QFI as a function of $h_{a}$ for various cycling periods when the periodic-field is in resonance with the binary quench, namely $T_{ac}{=}T$. Obviously, the QFI behaves differently in each phase. While in the DTC phase, one observes a plateau that uniformly increases with $n$, in the non-DTC region, non-trivial oscillations and suppression of the QFI are evident. The QFI clearly peaks at the same transition point denoted by $h_{a}^{\max}$ (dashed line) in Fig. 1 (b). The peaks become sharper as $n$ increases. Fig. 2 (b) depicts the dynamics of the QFI in three various regions, deep inside the DTC ($h_{a}{=}10^{-5}$), at the transition point ($h_{a}{=}h_{a}^{\max}$) and deep inside the non-DTC ($h_{a}{=}1$). The results show the square growth of the QFI, namely $\mathcal{F}_{Q}{\propto}n^{2}$ in the DTC phase. The unlimited quantum correlations at the transition point result in a faster growth of the QFI over early multiple cycles. However, at later times, one has $\mathcal{F}_{Q}{\propto}n^{2}$. In the Non-DTC region, the origin of QFI growth in later time is the contribution of several Floquet states that have considerable overlap with $|\psi(0)\rangle$ [111]. To extract the scaling behavior of the QFI, in Fig. 2 (c) we plot the obtained QFI as a function of $h_{a}$ after $n{=}10$ cycling period when the size of the chains varies. In both DTC and the transition point, the finite-size effect is apparent. By enlarging the chains, the peak of the QFI as the indicator of the phase transition skew toward smaller values of $h_{a}$ with a rate of $h_{a}^{\max}{\propto}L^{-1.0916}$, as shown in Fig. 2 (d). The result aligns with the analytical requirements of $h_{a}{\ll}\sqrt{2}\pi^{2}{/}L$ to achieve the ultimate precision. Although this result indicates degradation of the DTC order in the thermodynamic limit due to the presence of the periodic-field, in practice, it facilitates ultraprecise sensing of infinitesimal periodic-fields. In Fig. 2 (d), we show the values of the QFI deep inside the DTC phase ($h_{a}{=}10^{-5}$) and at the transition point ($h_{a}{=}h_{a}^{\max}$) when the length of the chains varies. The numerical results (markers) are well described by fitting the function $\mathcal{F}_{Q}{\propto}L^{\beta}$ with $\beta{=}4.0902$ and $\beta{=}2.5887$ for a sensor working inside the DTC order and at the transition point, respectively. This evidence indicates that leveraging a DTC system that is initialized in a product state $|\psi(0)\rangle$ results in a significant quantum enhancement in the sensing of the periodic-field.
@@ V.1 Off-resonance effect
Figure 3: (a) The imbalance as a function of $h_{a}$ for different values of the detuning $\delta f$ over different stroboscopic times. (b) The QFI versus the detuning $\delta f$ for various values of the $h_{a}$ and $n$. In both panels $L{=}7$. (c) The QFI in systems with different sizes as a function of $n$ when $h_{a}{=}10^{-2}$ and $\delta f{=}10^{-2}$. (d) The QFI as a function of $L$ when the sensor is in the DTC phase and at the transition point and $\delta f{=}10^{-2}$. The numerical results are well described by the fitting function $\mathcal{F}_{Q}{\propto}L^{\beta}$ with $\beta{=}4.6658$ and $\beta{=}2.2014$ for a sensor deep inside the DTC phase and at the transition point, respectively. In all the panels we set $\varepsilon{=}0.1$.
@@
While far off-resonance periodic-field has a negligible impact on the dynamic of the system, slight deviation denoted by $f_{ac}/f{=}1{+}\delta f$ with $\delta f$ as the frequency offset may affect the lifetime of the DTC order and the performance of the probe. In this case, one obtains $\Theta_{n}{=}\frac{(-1)^{n}h_{a}}{\pi(1+\delta f)}\big{(}\cos(n\pi\delta f){+}\cos((n{-}1)\pi\delta f)\big{)}$ as the accumulated phase in each period cycle. In Fig. 3(a) we present the imbalance as a function of $h_{a}$ over multiple period cycles when $\delta f$ varies. The response of $\mathcal{I}(nT)$ in the DTC region reveals the robustness of the system to small offset values. Fig. 3 (b) plots the QFI as a function of $\delta f$ over multiple period cycles when the probe works inside the DTC order. Clearly, the maximum values for the QFI happen for a resonant periodic-field. By increasing $\delta f$, the unfavorable off-resonance correlations affect the performance of the probe. Note that $\Theta_{n}{\rightarrow}0$ whenever $n\delta f{\rightarrow}0.5$. This affects the growth of the accumulation phase and therefore the value of the QFI as can be seen in Fig. 3 (c) where we plot the dynamics of the QFI in systems of various lengths under a periodic-field with magnitude $h_{a}{=}10^{-2}$ and offset $\delta f{=}10^{-2}$. All the accumulated phase in the first time window, namely for $\delta fn{<}0.5$, will be erased in the second time window, namely for $\delta fn{>}0.5$. This introduces an optimal sensing time as $n^{*}{=}1/2\delta f$, giving the peaks of the QFI. Interestingly, sensing in $n{\leqslant}n^{*}$ kept the scaling behavior of the QFI intact, as evidenced by Fig. 3 (d), where we plot the obtained $\mathcal{F}_{Q}$ for different system sizes when the probe works inside the DTC and at the transition point with offset $\delta f{=}10^{-2}$. Clearly, the obtained scaling $\beta$ is in line with our results for the resonance case, indicating the attainability of the quantum-enhanced sensitivity.
So far, we have focused on the scenario where the periodic-field is detected exclusively by subsystem $a$. However, in practice, the close distance between chains $a$ and $b$ introduces crosstalk, causing $b$ also to sense a portion of the periodic-field. When initializing the subsystems in polarized states, but with opposite total magnetization in the $z$-direction, the crosstalk influences the accumulated phase and, consequently, the value of the QFI.
@@ V.2 Cross-talk effect
Figure 4: (a) The QFI as a function of $h_{a}$ over various stereoscopic times, when $\eta{=}10^{-1}$. (b) The QFI as a function of $\eta$ when $n{=50}$. In both plots we set $L{=}7$ and $\varepsilon{=}0.1$.
@@
for $\mu{=}a,b$ in which $\eta{\in}(0,1)$. As evidenced by our results in Fig. 4 (a), where the QFI, as a function of $h_{a}$, has been plotted at different stroboscopic times when $\eta{=}10^{-1}$, the crosstalk can not affect the DTC order and the phase transition is driven by $h_{a}$. The distinct behavior of the QFI in both DTC and non-DTC regions, along with the enhanced peaks at the phase transition point, is evident. For completeness, in Fig. 4 (b) we present the QFI obtained as a function of $\eta$ when $h_{a}$ varies. The results are obtained after $n{=}50$ period cycling. Clearly, as the crosstalk parameter is increased, the values obtained for QFI decrease.
@@ V.3 Initialization effect
Having analyzed the effect of off-resonance and crosstalk, we now explore the robustness of our DTC probe to imperfections in the initialization. Regardless of the presence or absence of the periodic-field, as a perfect quench with $\varepsilon{=}0$ preserves $z$-polarization in each chain, the DTC order with arbitrary initialization is infinitely long-lived even in systems with finite sizes. For $\varepsilon{>}0$, the configuration basis states of $H_{I}$ exhibit varying levels of robustness to imperfections, with some remaining resilient at higher values of $\varepsilon$. In the previous section, we showed that initializing the system in state $|\psi(0)\rangle{=}|{\uparrow,\cdots,\uparrow}\rangle_{a}{\otimes}|{\downarrow,\cdots,\downarrow}\rangle_{b}$ provides persistence period-doubling oscillations of $\langle\mathcal{I}(nT)\rangle$ even for $\varepsilon{\neq}0$. Furthermore, significant overlap of this initial state with the ”Schrödinger cat” state $|\phi_{\pm}\rangle{=}\frac{1}{\sqrt{2}}(|\psi(0)\rangle{\pm}|\tilde{\psi}(0)\rangle)$ in the spectrum of the DTC has been identified as the main source of quantum-enhanced sensing. This obliges us to analyze the effect of perturbation in the initialization of the probe. For this, we consider a product state as
@@
$|\varsigma_{j}\rangle_{a}$ $={\rm cos(\alpha)}|\uparrow_{j}\rangle_{a}+{\rm sin(\alpha)}|\downarrow_{j}\rangle_{a}$ (14)
$|\varsigma_{j}\rangle_{b}$ $=-{\rm sin(\alpha)}|\uparrow_{j}\rangle_{b}+{\rm cos(\alpha)}|\downarrow_{j}\rangle_{b}$ (15)
with $\alpha{\in}[0,\pi/4]$, to avoid mixtures that yield $\langle\mathcal{I}(0)\rangle{=}0$. For the extreme values of $\alpha{=}0$, one can recover $|\psi(0)\rangle$. However, for other values of $\alpha$, a weighted mixture of different configurations is attainable. In Fig. 5 (a), we represent the normalized imbalance $\langle\mathcal{I}(nT)\rangle/\langle\mathcal{I}(0)\rangle$ over tens of period cycles when $\alpha$ varies. Clearly, the period-doubling oscillation of DTC remains unaffected by small imperfections during the initialization stage. To assess the performance of our DTC probe, in Fig. 5 (b), we plot the obtained QFI as a function of $h_{a}$ over different stroboscopic times when $\alpha{=}10^{-2}$. The results show that this level of perturbation in the initial state has no tangible impact on the performance of our DTC probe.
Figure 5: (a) The imbalance as a function of $n$ for different values of $\alpha$ when $h_{a}{=}10^{-2}$. (b) The QFI versus $h_{a}$ over different stroboscopic times when $\alpha{=}10^{-2}\pi$. In both panels $L{=}7$ and $\varepsilon{=}0.1$.
@@ V.4 Imperfect quench effect
After explaining the sharp phase transition driven by $h_{a}$, we will now investigate the melting of the DTC by increasing the imperfection $\varepsilon$. In Fig. 6(a), we plot the imbalance $\mathcal{I}(nT)$ after $n{=}50$ period cycles as a function of $h_{a}$ and $\varepsilon$. The phase diagram is entirely characterized by the parameters $h_{a}$ and $\varepsilon$. Increasing $\varepsilon$ reduces the DTC phase imbalance, but can improve the performance of the probe. As evident in Fig. 6(b), where we plot the obtained QFI as a function of $\varepsilon$ deep inside the DTC phase and at the transition point, by increasing the imperfection in the quench up to some value, the QFI grows. This growth in the DTC phase is more tangible than at the transition point. During the DTC phase, the dynamics is limited to a subspace spanned by $|\Phi_{\pm}\rangle$. Increasing $\varepsilon$ by engaging a larger portion of the Hilbert space encodes more information about $h_{a}$ in the quantum states. This finding contrasts sharply with typical sensors, where imperfections diminish performance.
@@
Figure 6: (a) The imbalance as a function of $h_{a}$ and $\varepsilon$ at $n{=}50$. (b) The QFI versus $\varepsilon$ for different values of the $h_{a}$ at $n{=}20$. In both panels $L{=}7$.
VI Experimental proposal
The methodology presented here is independent of the underlying physics, making it applicable to a broad spectrum of quantum and even classical systems. Here, we provide a proof-of-principle demonstration of our DTC sensor in ultracold atoms on an optical lattice, paving the way for realistic implementations.
Using the Jordan-Wigner transformation, one can easily map the Hamiltonian Eq. (4) to the spinless fermion model as

### v2

@@ IV DTC sensor
Fig. 2(a) presents the QFI as a function of $h_{a}$ for various cycling periods when the periodic-field is in resonance with the DTC, namely $T_{ac}{=}T$. Obviously, the QFI behaves differently in each phase. While in the DTC phase, one observes a plateau that uniformly increases with $n$, in the non-DTC region, non-trivial oscillations and suppression of the QFI are evident. The QFI clearly peaks at the same transition point denoted by $h_{a}^{\max}$ (dashed line) in Fig. 1(b). The peaks become sharper as $n$ increases. Fig. 2(b) depicts the dynamics of the QFI in three various regions, deep inside the DTC ($h_{a}{=}10^{-5}$), at the transition point ($h_{a}{=}h_{a}^{\max}$) and deep inside the non-DTC ($h_{a}{=}1$). The results show the square growth of the QFI, namely $\mathcal{F}_{Q}{\propto}n^{2}$ in the DTC phase. The unlimited quantum correlations at the transition point result in a faster growth of the QFI over early multiple cycles. However, at later times, one has $\mathcal{F}_{Q}{\propto}n^{2}$. In the non-DTC region, the origin of QFI growth in later time is the contribution of several Floquet states that have considerable overlap with $|\psi(0)\rangle$ [112]. To extract the scaling behavior of the QFI, in Fig. 2(c) we plot the obtained QFI as a function of $h_{a}$ after n=10n{=}10 cycling period when the size of the chains varies. In both DTC and the transition point, the finite-size effect is apparent. By enlarging the chains, the peak of the QFI as the indicator of the phase transition skew toward smaller values of $h_{a}$ with a rate of $h_{a}^{\max}{\propto}L^{-1.0916}$, as shown in Fig. 2(d). The result aligns with the analytical requirements of $h_{a}{\ll}\sqrt{2}\pi^{2}{/}L$ to achieve the ultimate precision. This means that, in the thermodynamic limit ($L\to\infty$), even an infinitesimal periodic field can destabilize the DTC order. While this implies fragility of the DTC phase in large systems due to the presence of the periodic-field, it simultaneously enables the system to act as an extremely sensitive detector, as vanishingly small field produces a pronounced dynamical response, effectively amplifying the system’s susceptibility to weak signals. This sharp transition in the dynamics can be exploited to detect the presence and properties of extremely weak periodic fields, making the system well-suited for high-precision metrology in practice. In Fig. 2(d), we show the values of the QFI deep inside the DTC phase ($h_{a}{=}10^{-5}$) and at the transition point ($h_{a}{=}h_{a}^{\max}$) when the length of the chains varies. The numerical results (markers) are well described by fitting the function $\mathcal{F}_{Q}{\propto}L^{\beta}$ with $\beta{=}4.0902$ and $\beta{=}2.5887$ for a sensor working inside the DTC order and at the transition point, respectively. This evidence indicates that leveraging a DTC system that is initialized in a product state $|\psi(0)\rangle$ results in a significant quantum enhancement in the sensing of the periodic-field, represented by
$\mathcal{F}_{Q}\propto L^{\beta}n^{\alpha}\quad\mathrm{with}\quad\beta{\simeq}4,\quad\alpha{\simeq}2.$ (14)
@@ V.1 Off-resonance effect
Figure 3: (a) The imbalance as a function of $h_{a}$ for different values of the detuning $\delta f$ over different stroboscopic times. (b) The QFI versus the detuning $\delta f$ for various values of the $h_{a}$ and $n$. In both panels L=7L{=}7. (c) The QFI in systems with different sizes as a function of $n$ when $h_{a}{=}10^{-2}$ and $\delta f{=}10^{-2}$. (d) The QFI as a function of $L$ when the sensor is in the DTC phase and at the transition point and $\delta f{=}10^{-2}$. The numerical results are well described by the fitting function $\mathcal{F}_{Q}{\propto}L^{\beta}$ with $\beta{=}4.6658$ and $\beta{=}2.2014$ for a sensor deep inside the DTC phase and at the transition point, respectively. In all the panels, we set $\varepsilon{=}0.1$.
@@
While far off-resonance periodic-field has a negligible impact on the dynamic of the system, slight deviation denoted by $f_{ac}/f{=}1{+}\delta f$ with $\delta f$ as the frequency offset of the DTC’s frequency may affect the lifetime of the DTC order and the performance of the probe. In this case, one can again approximate the sinusoidal periodic-field with a square-wave form of the same period and obtain the total accumulated phase through integrating over $h_{a}\sin(2\pi(1+\delta f)f\tau)$ over the two halves of each period $T$, namely ${(n-1)T},{(n-\frac{1}{2})T}]$ and ${(n-\frac{1}{2})T},{nT}]$. Therefor, the total accumulated phase in each period cycle $T$ is $\frac{(-1)^{n+1}h_{a}}{\pi(1+\delta f)}\big(\cos(n\pi\delta f){+}\cos((n{-}1)\pi\delta f)\big)$. In Fig. 3(a) we present the imbalance as a function of $h_{a}$ over multiple period cycles when $\delta f$ varies. The response of $\mathcal{I}(nT)$ in the DTC region reveals the robustness of the system to small offset values. Fig. 3(b) plots the QFI as a function of $\delta f$ over multiple period cycles when the probe works inside the DTC order. Clearly, the maximum values for the QFI happen for a resonant periodic-field. By increasing $\delta f$, the unfavorable off-resonance correlations affect the performance of the probe. Note that $\Theta_{n}{\rightarrow}0$ whenever $n\delta f{\rightarrow}0.5$. This affects the growth of the accumulation phase and therefore the value of the QFI as can be seen in Fig. 3(c) where we plot the dynamics of the QFI in systems of various lengths under a periodic-field with magnitude $h_{a}{=}10^{-2}$ and offset $\delta f{=}10^{-2}$. All the accumulated phase in the first time window, namely for $\delta fn{<}0.5$, will be erased in the second time window, namely for $\delta fn{>}0.5$. This introduces an optimal sensing time as $n^{*}{=}1/2\delta f$, giving the peaks of the QFI. Interestingly, sensing in $n{\leqslant}n^{*}$ kept the scaling behavior of the QFI intact, as evidenced by Fig. 3(d), where we plot the obtained $\mathcal{F}_{Q}$ for different system sizes when the probe works inside the DTC and at the transition point with offset $\delta f{=}10^{-2}$. Clearly, the obtained scaling $\beta$ is in line with our results for the resonance case, indicating the attainability of the quantum-enhanced sensitivity. So far, we have focused on the scenario where the periodic-field is detected exclusively by subsystem $a$. However, in practice, the close distance between chains $a$ and $b$ introduces crosstalk, causing $b$ also to sense a portion of the periodic-field. When initializing the subsystems in polarized states, but with opposite total magnetization in the $z$-direction, the crosstalk influences the accumulated phase and, consequently, the value of the QFI.
@@ V.2 Cross-talk effect
Figure 4: (a) The QFI as a function of $h_{a}$ over various stereoscopic times, when $\eta{=}10^{-1}$. (b) The QFI as a function of $\eta$ when n=50n{=50}. In both plots we set L=7L{=}7 and $\varepsilon{=}0.1$.
@@
for $\mu{=}a,b$ in which $\eta{\in}(0,1)$. As evidenced by our results in Fig. 4(a), where the QFI, as a function of $h_{a}$, has been plotted at different stroboscopic times when $\eta{=}10^{-1}$, the crosstalk can not affect the DTC order and the phase transition is driven by $h_{a}$. The distinct behavior of the QFI in both DTC and non-DTC regions, along with the enhanced peaks at the phase transition point, is evident. For completeness, in Fig. 4(b) we present the QFI obtained as a function of $\eta$ when $h_{a}$ varies. The results are obtained after n=50n{=}50 period cycling. Clearly, as the crosstalk parameter is increased, the values obtained for QFI decrease.
@@ V.3 Initialization effect
Figure 5: (a) The imbalance as a function of $n$ for different values of $\vartheta$ when $h_{a}{=}10^{-2}$. (b) The QFI versus $h_{a}$ over different stroboscopic times when $\vartheta{=}10^{-2}\pi$. In both panels L=7L{=}7 and $\varepsilon{=}0.1$.
Having analyzed the effect of off-resonance and crosstalk, we now explore the robustness of our DTC probe to imperfections in the initialization. Regardless of the presence or absence of the periodic-field, as a perfect quench with $\varepsilon{=}0$ preserves total $z$-polarization in the system, the DTC order with arbitrary initialization is infinitely long-lived even in systems with finite sizes. For $\varepsilon{>}0$, the configuration basis states of $H_{I}$ exhibit varying levels of robustness to imperfections, with some remaining resilient at higher values of $\varepsilon$. In the previous section, we showed that initializing the system in state $|\psi(0)\rangle{=}|{\uparrow,\cdots,\uparrow}\rangle_{a}{\otimes}|{\downarrow,\cdots,\downarrow}\rangle_{b}$ provides persistence period-doubling oscillations of $\langle\mathcal{I}(nT)\rangle$ even for $\varepsilon{\neq}0$. Furthermore, significant overlap of this initial state with the ”Schrödinger cat” state $|\Phi_{\pm}\rangle{=}\frac{1}{\sqrt{2}}(|\psi(0)\rangle{\pm}|\tilde{\psi}(0)\rangle)$ in the spectrum of the DTC has been identified as the main source of quantum-enhanced sensing. This obliges us to analyze the effect of perturbation in the initialization of the probe. For this, we consider a product state as
@@
$|\varsigma_{j}\rangle_{a}$ $={\rm cos(\vartheta)}|\uparrow_{j}\rangle_{a}+{\rm sin(\vartheta)}|\downarrow_{j}\rangle_{a}$ (17)
$|\varsigma_{j}\rangle_{b}$ $=-{\rm sin(\vartheta)}|\uparrow_{j}\rangle_{b}+{\rm cos(\vartheta)}|\downarrow_{j}\rangle_{b}$ (18)
with $\vartheta{\in}[0,\pi/4]$, to avoid mixtures that yield $\langle\mathcal{I}(0)\rangle{=}0$. For the extreme values of $\vartheta{=}0$, one can recover $|\psi(0)\rangle$. However, for other values of $\vartheta$, a weighted mixture of different configurations is attainable. In Fig. 5(a), we represent the normalized imbalance $\langle\mathcal{I}(nT)\rangle/\langle\mathcal{I}(0)\rangle$ over tens of period cycles when $\vartheta$ varies. Clearly, the period-doubling oscillation of DTC remains unaffected by small imperfections during the initialization stage. To assess the performance of our DTC probe, in Fig. 5(b), we plot the obtained QFI as a function of $h_{a}$ over different stroboscopic times when $\vartheta{=}10^{-2}\pi$. The results show that this level of perturbation in the initial state has no tangible impact on the performance of our DTC probe.
@@ V.4 Imperfect quench effect
Figure 6: (a) The imbalance as a function of $h_{a}$ and $\varepsilon$ at n=50n{=}50. (b) The QFI versus $\varepsilon$ for different values of the $h_{a}$ at n=20n{=}20. In both panels L=7L{=}7.
After explaining the sharp phase transition driven by $h_{a}$, we will now investigate the melting of the DTC by increasing the imperfection $\varepsilon$. In Fig. 6(a), we plot the imbalance $\mathcal{I}(nT)$ after n=50n{=}50 period cycles as a function of $h_{a}$ and $\varepsilon$. The phase diagram is entirely characterized by the parameters $h_{a}$ and $\varepsilon$. Increasing $\varepsilon$ reduces the DTC phase imbalance, but can improve the performance of the probe. As evident in Fig. 6(b), where we plot the obtained QFI as a function of $\varepsilon$ deep inside the DTC phase and at the transition point, by increasing the imperfection in the quench up to some value, the QFI grows. This growth in the DTC phase is more tangible than at the transition point. During the DTC phase, the dynamic is limited to a subspace spanned by $|\Phi_{\pm}\rangle$. Increasing $\varepsilon$ by engaging a larger portion of the Hilbert space encodes more information about $h_{a}$ in the quantum states. This finding contrasts sharply with typical sensors, where imperfections diminish performance.
VI Feasibility of quantum-enhanced sensitivity
Figure 7: (a) Time-averaged CFI over N=50N{=}50 as a function of $L$ in DTC phase, namely $h_{a}{=}10^{-5}$ obtained using computational-basis (total magnetization) measurement, denoted by $\overline{\mathcal{F}}_{C}$ ($\overline{\mathcal{F}}^{S^{z}}_{C}$) for the probes in resonance ($\delta f{=}0$) and out of the resonance ($\delta f{=}0.01$) with the periodic-field. The markers show numerical results, and the solid lines represent the best-fit function as $\overline{\mathcal{F}}_{C}{\propto}L^{\beta}$ with reported $\beta$’s. Inset is $\overline{\mathcal{F}}_{C}$ and time-averaged QFI, $\overline{\mathcal{F}}_{Q}$, over N=100N{=}100 as a function of $h_{a}$ in a chain of size L=7L{=}7 and $\delta f{=}0$. (b) Point averaged QFI and CFI versus $n_{av}$ in a system of size L=7L{=}7 that operates in the DTC phase, namely $h_{a}{=}10^{-5}$. The markers are the numerical results, and the solid lines are the best fitting function as $\overline{\mathcal{F}}_{Q,C}{\propto}n_{av}^{\alpha}$ with reported alphaalpha’s. All the results are obtained for $\varepsilon{=}0.1$.
Saturating the quantum Cramér-Rao bound requires measuring the system using optimal measurements, which are generally complex and may even depend on the unknown parameter $h_{a}$. Therefore, finding an experimentally feasible set of measurements that results in $\mathcal{F}_{C}$ close to $\mathcal{F}_{Q}$ is highly desirable. Interestingly, in our sensing scheme, simple measurements such as computational-basis measurement, described by $\{\Pi_{\mathbf{z}}{=}\left|\mathbf{z}\right\rangle\left\langle\mathbf{z}\right|\}$ where in $\left|\mathbf{z}\right\rangle$ represents $2^{2L}$ different computational basis, or collective-spin observable of each chains described by $S^{\mu z}{=}\sum_{j=1}^{L}\sigma^{\mu z}_{j}$, can follow the scaling behavior of the QFI, concerning both time and size. Here, we calculate the CFI for the computational-basis measurement using $p_{\mathbf{z}}{=}|\langle\mathbf{z}|\psi(n)\rangle|^{2}$ and for the collective-spin observable using $p{=}\langle\psi(n)|S^{az}|\psi(n)\rangle$. The results are denoted as $\overline{\mathcal{F}}_{C}$ and $\overline{\mathcal{F}}^{S^{z}}_{C}$, respectively. The inset of Fig. 7(a) provides time-averaged CFI, defined as $\overline{\mathcal{F}}_{C}{=}\frac{1}{N}\sum_{n=1}^{N}\mathcal{F}_{C}$, and the time-averaged QFI as $\overline{\mathcal{F}}_{Q}{=}\frac{1}{N}\sum_{n=1}^{N}\mathcal{F}_{Q}$. Here, $N$ is the total available time for sensing. Similar to the QFI, the time-averaged CFI shows distinct behaviors in both the DTC phase and the non-DTC region. In addition, it peaks sharply at the transition point $h_{a}^{\max}$ and highly resembles the QFI. In Fig. 7(a), we show the scaling behavior of $\overline{\mathcal{F}}_{C}$, when the probe operates in the DTC phase, namely $h_{a}{=}10^{-5}$. Our results reveal that $\overline{\mathcal{F}}_{C}$ follows the scaling behavior as $\overline{\mathcal{F}}_{C}{\propto}L^{\beta}$ with obtained $\beta{\simeq}3.98$ through fitting process. Same scaling behavior is observed for $\overline{\mathcal{F}}^{S^{z}}_{C}{=}\frac{1}{N}\sum_{n=1}^{N}\mathcal{F}^{S^{z}}_{C}$, as can be seen from Fig. 7(a). In this case, we obtain $\beta{\simeq}3.99$. To extract the scaling behavior concerning time, we divide the total sensing time $N$ into $K$ smaller time intervals with equal space $n_{i+1}{-}n_{i}{=}\Delta n$ and average both QFI and CFI over these intervals. Fig. 7(b) represents the results as a function of $n_{av}{=}\sum_{i=1}^{K}(i\Delta n)$. While the markers are numerical results, the solid lines are the best fitting function as $\overline{\mathcal{F}}_{Q},\overline{\mathcal{F}}_{C},\overline{\mathcal{F}}^{S^{z}}_{C}{\propto}n_{av}^{\alpha}$ with $\alpha{=}2.014$ for the QFI, $\alpha{=}2.0736$ for the CFI using computational-basis measurement, and $\alpha{=}2.0942$ for the CFI using collective-spin measurement. These numerical analyses show that the simple measurement setups in our DTC probe can capture the quantum-enhanced sensitivity concerning both time and the size of the probe, represented by
$\overline{\mathcal{F}}_{C}\propto L^{\beta}n^{\alpha}_{av}\quad\mathrm{with}\quad\beta{\simeq}4,\quad\alpha{\simeq}2.$ (19)
This scaling behavior remains unchanged regardless of the frequency offset of the field. In Fig. 7(a), we plot the CFI using the computational-basis measurements as a function of $L$ for sensing the periodic-field with frequency offset $\delta f{=}0.01$. The numerical results (markers) are properly described by the fitting function $\overline{\mathcal{F}}_{C}\propto L^{\beta}$ with $\beta{\simeq}3.89$.
VII Effect of Decoherence
Regardless of how much effort is put into isolation, quantum systems inevitably interact with their environment. These interactions can cause decoherence, dissipation, and energy exchange. Introducing an external environment, without fine-tuning, can disrupt subharmonic oscillations and thermalize the system [119, 112, 120]. In this section, we analyze the effect of local dephasing on the sensing capability of the DTC probe. Considering Markovian diphasing noise, the system’s density matrix evolves according to the time-dependent Lindblad equation as
$\partial_{t}\rho=\begin{cases}-i[H_{a}+H_{b}+\Theta_{n}G^{az},\rho]+\Gamma\big(\mathcal{L}^{a}[\rho]+\mathcal{L}^{b}[\rho]\big)&(n-1)T<t\leqslant(n-\frac{1}{2})T\\ -i[H_{I}+\Theta_{n}G^{az},\rho]+\Gamma\big(\mathcal{L}^{a}[\rho]+\mathcal{L}^{b}[\rho]\big)&(n-\frac{1}{2})T<t\leqslant nT\end{cases}$
here, $\mathcal{L}^{\mu}[\rho]{=}\sum_{j=1}^{L}\big(L_{j}^{\mu\dagger}\rho L^{\mu}_{j}-\frac{1}{2}L_{j}^{\mu\dagger}L^{\mu}_{j}\rho-\frac{1}{2}\rho L_{j}^{\mu\dagger}L^{\mu}_{j}\big)$ represents the Liouvillian super-operator resulting from the coupling to the environment. $L^{\mu}_{j}{=}\sigma^{\mu z}_{j}$ is the Lindblad operator representing the dephasing process with rate $\Gamma$ in chain $\mu{=}a,b$. To extract the scaling behavior with respect to time, in Fig. 8, we depict the point-averaged QFI and CFI as functions of $n_{av}$ in systems of size (a) L=3L{=}3, and (b) L=4L{=}4, when the probe operates in the DTC phase, namely $h_{a}{=}10^{-5}$, and exposed to the dephasing noise with rate $\Gamma{=}10^{-3}$. The behavior of numerical results (marker) can be described by the fitting function $\overline{\mathcal{F}}_{Q,C}\propto n_{av}^{\alpha}$ with the attached $\alpha$’s to the plots. Two main observations are in order. First, even with dephasing noise, a quantum enhancement in time, namely $\alpha>1$, is feasible. Second, as shown in Fig. 8, increasing the system size leads to a higher $\alpha$’s, resulting in a greater quantum enhancement.
Figure 8: Point-averaged QFI and CFI versus $n_{av}$ in a system of size (a) L=3L{=}3 and (b) L=4L{=}4 that operates in the DTC phase, namely $h_{a}{=}10^{-5}$, in the presence of local dephasing with rate $\Gamma{=}10^{-3}$. The markers are the numerical results, and the solid lines are the best fitting function as $\overline{\mathcal{F}}_{Q,C}\propto n_{av}^{\alpha}$. All the results are obtained for $\varepsilon{=}0.1$.
VIII Experimental proposals
The methodology presented here is independent of the underlying physics, making it applicable to a broad spectrum of quantum and even classical systems. Here, we provide a proof-of-principle demonstration of our DTC sensor in ultracold atoms on an optical lattice, which offers a promising platform for experimental realization of the proposed DTC due to its broad controllability and flexibility. Specifically, the precisely tunable height and width of the lasers used to create the optical trapping potential enable the demanding quench required by our protocol. A detailed discussion on implementing this DTC using dipolar gases in an optical lattice can be found in Ref. [73]. Using the Jordan-Wigner transformation, one can easily map the Hamiltonian Eq. (4) to the spinless fermion model as